OpenStudy (christina166):

A circle is centered at (−4, −6) and has a radius of 3. Which of the following is the equation for this circle? (x + 4)2 + (y + 6)2 = 3 (x − 4)2 + (y − 6)2 = 9 (x − 4)2 + (y − 6)2 = 3 (x + 4)2 + (y + 6)2 = 9

OpenStudy (christina166):

@pooja195 May you please help.

jagr2713 (jagr2713):

What do we do first?

OpenStudy (marcelie):

hmm.. you can plugin the points onto ur equation

OpenStudy (christina166):

I really don't get it :/

OpenStudy (bobo-i-bo):

Hmmm, do you understand/know the equation for a circle of radius one, centered at the point (0,0)?

OpenStudy (christina166):

Not really, I'm sorry I feel stupid.

OpenStudy (mathmale):

Most basic equation of a circle is the one centered at the origin, (0,0):\[x^2 + y^2 = r^2\]

OpenStudy (mathmale):

If, on the other hand, your circle is centered at (h,k), the basic equation for that circle of radius r is\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (christina166):

Oh okay.

OpenStudy (mathmale):

Now you're given the coordinates of the center of the circle. They are (-4,-6), and the radius is 3.

OpenStudy (mathmale):

Substitude -4 for x, -6 for k and 3 for r into the most recent equation I gave you.

OpenStudy (christina166):

Oh, thank you!

OpenStudy (mathmale):

Next time you feel stuck on a problem like this one, look up the formula(s) on the 'Net. I would have looked up "equation of a circle." try it. What a great tool the 'Net can be! You're very welcome, Christina. See you again.

OpenStudy (christina166):

Okay, is it okay if you help me with the last question I was stuck with? @mathmale

OpenStudy (mathmale):

Sure. Please post it here.